CMS FED Testing. Update 28-06-2002 M. Noy & J. Leaver Imperial College Silicon Group. DAC Evaluation board. LVDS. UTP. LVTTL. SEQSI. UTP. VME. Opto-Tx. I2C control. I2C master. Optical Fibre. 8 bit oscilloscope. Coax. Opto-Rx. GPIB. PC. CMS FED Testing. Currently:.
Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.
M. Noy & J. Leaver
8 bit oscilloscope
CMS FED Testing
Line driver + (optional) level shift
Picture of the DAC evaluation board
LVTTL data in, from level converter
Picture of the key link components
UTP from DAC/line driver
Single ended output to scope
CMS FED Testing
DAC Evaluation Board
Signal before the link
200mV (V+-V- =400mV) differential signal with no offset.
higher bandwidth ringing and faster rise/fall time.
Time scale is relative to the scope trigger point on all plots.
Signal after the link
Single ended, with offset.
No ringing but slower
Signal noise/jitter after the link
Note:no scale on the width of the line. This is an impression of the infinite persistence scope trace => spread unknown.
(Measurements are real).
Rise time, 10% to 90% of full scale.
Fall time 90% to 10% of full scale
Linearity: link is being operated in the linear region of the Tx/Rx
Settings: x0, x1, x2, x3, x4, x5=0,0,0,0,1 (recommended by CERN)
Sample APV25 pair of multiplexed frames with simulated 1 MIP signal
Multiplexed APV25 header with zero pipeline address
2 error bits
address (16 bits)
2x12x25ns bits =
6 start bits
Zoom in of the 1 MIP signal upon its pedestal
We attached a heating element and a thermocouple to the laser package and used the following PID equation to stabilise the temperature through a feedback loop.
W = P [ (Ts - T0) + D d (Ts- T0) /dt + I (Ts - T0)dt]
The temperature was varied between 30ºC and 40ºC in ~1ºC steps. The output of the Rx was recorded after a stabilisation time, for some measurement time.
Laser threshold bias current behaves like
And (after a few lines and other things!)
Vout -Reff G Rx l Ith T/T0
Typical parameter values yield an expected (@ 34.1°C)
Vout/T -90.8 mV/°C
Temperature Measurement Results
Approximated with a linear fit V=mT+C
Where m = -(89.8 1.8) mV/°C
and C = (4137 62) mV
Good agreement with expected value (of 90.8mV/°C), but some of the parameters are loosely defined
Temperature Measurement Errors
Temperature stabilisation is good, with random fluctuations of the order 0.02°C. There is some unknown systematic error, that does not exceed 0.44°C.
Have statistics of 150x500 voltage points and 150 temperature points per temperature setting.
Statistical errors are too small to account for the largest random deviation, probably spurious.
We could repeat the whole measurement again using smaller T steps, but probably won’t due to time constraints.
Have a complete working single fibre, possible to drive 4 with identical signals using the current Opto-Tx.
Possible to obtain a further 2 of the 4 channel prototypes from CERN complete 12 channels could in principle be driven with identical signals.
Dependence of laser operation on temperature is now better understood, and fine temperature control is possible. We feel confident that a system such as the one we have will allow sufficient temperature stability for the fed testing needs.
Work is in progress to produce an application specific version of the SEQSI
longer RAM pipeline
clean/synchronous stop from VME
possible stepping through
Have 1 (untested) Opto-Rx emulator to drive the analogue stage of the FED directly over copper (I.e. eliminating the optical link)
Verify DAC Linearity: Summer student(?)
More thought into a test vectors and their comparison with the FED output